The total cost of five ice creams and four smoothies is 9.20. The total cost of three ice creams and six smoothies id 8.40
5x + 4y =9.20
3x + 6y =8.40


By solving these equations find the cost of an ice cream and the cost of a smoothie.

Answer by checkley79(3341)

5x + 4y =9.20 MULTIPLY BY 3
3x + 6y =8.40 MULTIPLY BY-2 & ADD
15X+12Y=27.60
-6X-12Y=-16.80
----------------
9X=10.80
X=10.80/9
X=1.20 FOR EACH ICE CREAM
5*1.20+4Y=9.20
6.00+4Y=9.20
4Y=9.20-6.00
4Y=3.20
Y=3.20/4
Y=.80 FOR EACH SMOOTHIE
PROOF:
3*1.20+6*.80=8.40
3.60+4.80=8.40
8.40=8.40